OpenStudy (anonymous):

If f(x) = |x – 5| + 2, find f(3)

OpenStudy (anonymous):

what's |3-5| = ?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

no

OpenStudy (anonymous):

absolute value of -2 = ?

OpenStudy (anonymous):

2

OpenStudy (anonymous):

add 2 to that = ?

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

why 2+2 = 0

OpenStudy (anonymous):

add 2 to 2

OpenStudy (anonymous):

oh 4

OpenStudy (anonymous):

thats ur answer do u understand absolute values |...| ?

OpenStudy (anonymous):

they always generate a positive value, never a negative, ok?

OpenStudy (anonymous):

alright, thanks. do you know how to do these? What is the turning point of the graph of f(x) = |x + 3| – 2? /:

OpenStudy (anonymous):

 will be something to the above, you need to find where it turns

OpenStudy (anonymous):

how do you find that?

OpenStudy (anonymous):

when f(x) = 0, what's x and where is the y intercept?

OpenStudy (anonymous):

-3 is x and y is -2?

OpenStudy (anonymous):

or... 3 and 2?

OpenStudy (anonymous):

f(x) = 0 |x+3|-2 = 0 |x+3| = 2 x = -1, because -1+3 = 2 and x = -5 because -5+3 = -2 but we have absolute value around it so answer = 2

OpenStudy (anonymous):

okay,so is the answer (2,-3) or (-3,2)?

OpenStudy (anonymous):

to find the u intercept, put x = 0, what f(x)

OpenStudy (anonymous):

sorry Y not u

OpenStudy (anonymous):

so it's (2, -3)?

OpenStudy (anonymous):

y intercept is 1

OpenStudy (anonymous):

well 1 isn't an answer choice so......

OpenStudy (anonymous):

no I know ... we found the y intercept only

OpenStudy (anonymous):

ok cool i have to turn this in in 8 minutes sooooooo what do i do now.

OpenStudy (anonymous):

i know it's 2, -3 or -3, 2 so which order.

OpenStudy (anonymous):

-3,2 then

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

what answers do you have, all options? is it multiple choice

OpenStudy (dan815):

for absolute values u are always taking the magnitude, if its a negative number u still take it as a positive number

OpenStudy (anonymous):

best if you post the second question as a standalone question so others can help too

OpenStudy (anonymous):

is there an answer of -3,-2 ?

OpenStudy (anonymous):